You have x grams of Na2Cr2O7
How many grams of CrCl3 will be formed if the Na2Cr2O7 undergoes the reaction below?

Na2Cr2O7+3Zn+14HCl --> 2CrCl3+3ZnCl+2NaCl+7H2O
1
Expert's answer
2020-08-27T05:04:39-0400

Na2Cr2O7 + 3Zn + 14HCl -> 2CrCl3 + 3ZnCl2 + 2NaCl + 7H2O

find the amount of substance (n):


"\frac{x}{1}=\frac{y}{2}"

y=2x


"m=M\times n=" 158.3550\times2x=316.71x

M(ClCr3)=158.3550